OpenStudy (anonymous):

if y= 4, what is the answer to the following expression 3/y+2y/4

OpenStudy (anonymous):

plug in 4 everywhere there is a y...gives you 3/4+2=11/4

OpenStudy (nottim):

Just wondering, does the / represent divide?

OpenStudy (anonymous):

yea...usually representative of a fraction w/o using the eq. editor

OpenStudy (nottim):

but then its 3/11 and then divided 4? That's not correct form.

OpenStudy (anonymous):

3/11?

OpenStudy (nottim):

plug in 4 everywhere there is a y...gives you 3/4+2=11/4. "3/4+2=11"

OpenStudy (nottim):

IS mekax even here anymore?

OpenStudy (anonymous):

how so? 3 + 8 = 11 - - - 4 4 4

OpenStudy (anonymous):

yes im here

OpenStudy (nottim):

Try using the Equation or Draw Thing tmoney. Im not understanding you.

OpenStudy (anonymous):

ok

OpenStudy (nottim):

Thank you.

OpenStudy (anonymous):